Enlarge the footprint by 5 feet for accuracy and slope, then determine the volume of topsoil to be removed and stockpiled, if the thickness 9 inches. If no soil is to be stockpiled, and the swell is 10 percent, how many truck loads will it take to cart away from the site using 7 cy trucks? cubic yards

100 120 240 120 180 2 VT
Answer #1

We have,

8+5 10 5 5 10 12 5 12 12 12 5 5 18 10+5

(Gaps are 5 feet in measurement)

Finding volume of topsoil to be removed and stockpiled:

There is a 5 feet gap near foundation sides, for workers to stand. Then the area of excavation is,

10 + 5 + 5 = 20 ft

12 + 5 = 17 ft

20 \times 17 = 340 ft2

8 + 5 = 13 ft

12 + 5 + 5 = 22 ft

13 \times 22 = 286 ft2

10 + 5 = 15 ft

12 + 5 = 17 ft

15 \times 17 = 255 ft2

340 ft2 + 286 ft2 + 255 ft2 = 881 ft2

The area of excavation is, 881 ft2

Given, thickness 9 inch = (9/12) ft

Now,

881 ft2 (9/12) ft = 660.75 ft3

Converting into cubic yards (yd3)

660.75 \times (1/27) = 24.47 yd3

The volume of topsoil to be removed and stockpiled is, 24.47 cubic yards.

Finding the volume of soil after the swelling of 10% = 0.10:

24.47 + 24.47(0.10) = 26.92 cubic yards

Each truck contains 7 cubic yards of soil, then total number of trucks required to carry away the soil is,

26.92/7 = 3.85 ~ 4

4 trucks required to carry away 26.92 cubic yards soil.